# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@ IV Scaling and fidelity for heralded schemes
where we include a superscript to denote emitter $(e)$ and auxiliary spin $(s)$ qubits, and we have used the same coherence time $\tau$ for the emitter and auxiliary spins in the system.
For the HoS scheme, we assume a standard SPDC source to generate the required entangled photon pairs [50]. The multi-pair emission of such a source introduces an additional infidelity in the entanglement swapping procedure. We denote $P_{t}$ as the overall probability that a true Bell state is heralded and measured after constructing the graph. In general, $P_{t}<P_{s}$ due to false heralding. A primary source of false heralding is when the emitter photon is not collected but the SPDC source produces two or more photon pairs that lead to a detection pattern, falsely signaling a successful swap. False heralding due to two signal photons is a well known and ubiquitous problem in entanglement swapping with photon pair sources based on nonlinear optics [38]. We can write an expression for the fidelity of the entanglement swapping procedure, $P_{t}/P_{s}$, using the known statistical distribution of entangled photon pairs produced via SPDC. Here, we assume photon number-resolved detection and perfect collection and detection of the signal photon once produced. This is because loss on that channel can generally be minimal and primarily affects the rate rather than the swapping fidelity (see Appendix D for the full calculation, including further experimental factors not discussed here). Furthermore, the intention here is to show that even quite low efficiency collection of the emitter photon can allow the generation of graph states. The probability of producing two entangled photon pairs is proportional to the square of the probability of producing a single entangled pair, and is controllable via the pump power of the SPDC (parameterized here by the quantity $\xi$ [38]), thus introducing a tradeoff between rate and fidelity. The fidelity, $F_{\rm{swap}}=P_{t}/P_{s}$, is given by
$P_{t}(\eta_{e},\xi)=\eta_{e}(1-\xi)^{2}\xi,$ (2a)
$P_{s}(\eta_{e},\xi)=\left[\eta_{e}\xi+(1-\eta_{e})\xi^{2}\right](1-\xi)^{2},$ (2b)
$F_{\rm{swap}}(\eta_{e},\xi)=\frac{\eta_{e}}{\eta_{e}+(1-\eta_{e})\xi}.$ (2c)
In the absence of decoherence of the spins, we could simply work at $\xi\ll 1$ (very low pair generation rate) to reduce the false heralding and increase the overall fidelity. However, this slows the rate of the graph production and introduces infidelity due to decoherence of the auxiliary spin(s). Given a set of experimental parameters, the rate of pair production can be optimized by tuning $\xi$ to maximize the overall fidelity of the desired graph state, balancing the infidelity due to false heralds against the decoherence.
We present a simple example comparing these additional sources of infidelity to the limits set by spin-spin entangling gates and initial entanglement fidelities, for the case of constructing a graph state on a single auxiliary spin. Figure 4 depicts these estimates over a few regimes. The combined effect of false heralds and decoherence (blue) set the fidelity as
$F(n_{p},\eta_{e},\xi)$ $=F_{\text{swap}}(\eta_{e},\xi)^{n_{p}}F_{D}^{(e)}(n_{p})$ (3)
×$\quad\quad\quad\quad\times\langle F_{D}^{(s)}(n_{p},P_{s}(\eta_{e},\xi))\rangle,$
which we optimize over $\xi$, for fixed $n_{p}$ and $\eta_{e}$. When only limited by decoherence (yellow), we can employ Equation (3), except now fixing $F_{\text{swap}}(\eta_{e},\xi)=1$ and removing the need to optimize the fidelity. These estimates are applicable when constructing large star graph states—the local-unitary equivalent to a Greenberger-Horne-Zeilinger (GHZ) state—and 1D cluster states.
V Realizing secure two-party computation
@@
As one application of our HoD scheme, we describe a method for securely computing an arbitrary Boolean function $f:\{0,1\}^{\times n}\to\{0,1\}$ of either two parties or a restricted class of multi-party functions. As a special type of MBQC, our protocol, $\mathcal{P}$, performs the calculation through a sequence of measurements on distributed graph states. Only Pauli measurements are needed, so $\mathcal{P}$ can be implemented using HoD and the generation of the requisite resource state happens in parallel with the computation without any need for photonic memory. Furthermore, the size of each graph state required for $\mathcal{P}$ is fixed at $12$ photons, regardless of the number of parties or the size of their inputs, which should be possible with high fidelity on current state-of-the-art hardware.
@@
We propose an unconditionally secure method for restricted MPC that includes all two-party computations and requires only two rounds of public communication regardless of the size of the computational input. Our scheme follows a well-known approach of decomposing an MPC into “offline” and “online” phases [54, 55, 56]. In the offline phase, some universal computational resource is distributed to all the parties. Crucially, this resource does not depend on the particular function being computed other than its input size. Then in the online phase this resource is used to compute some chosen function of the parties’ inputs. For example, in the classical setting one well-known computational resource is a special form of shared randomness known as “Beaver triples,” which can be used to efficiently compute logical AND gates in the online phase [57]. The problem of MPC then reduces to secure and efficient methods for distributing Beaver triples in the offline phase. In a similar spirit, our protocol involves distributing certain quantum graph states in the offline phase which then enable the computation of a logical AND in the online phase, through quantum measurement and classical post-processing. Beyond its relatively low communication costs, a significant advantage of our protocol is that the parties can, in principle, use self-testing methods to verify that some untrusted Source is faithfully distributing the correct graph state [58, 59], an ability that does not exist for classical shared randomness.
@@
Suppose that $N$ parties $\mathbf{P}_{1},\cdots,\mathbf{P}_{N}$ wish to compute some Boolean function $f(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$, where $\mathbf{x}_{k}$ is a string of bits representing the input for party $\mathbf{P}_{k}$. In addition to correctness, the evaluation of $f$ should be done securely such that the parties learn no more information about the individual $\mathbf{x}_{1},\cdots,\mathbf{x}_{N}$ beyond their own input and what is revealed in the function value $f(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$. To achieve this task, we propose a method of delegated computation in which a non-collaborating Referee, $\mathbf{R}$, is introduced to assist in the computation of $f(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$. To maintain privacy, $\mathbf{R}$ also should not learn any more information about the $\mathbf{x}_{k}$ beyond what is implied by the computed value $f(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$, nor does $\mathbf{R}$ reveal any more information to the other parties, other than what is necessary to evaluate $f$ securely.
@@
We utilize the fact that every Boolean function $f$ can be expressed in an algebraic normal form (ANF), which presents $f$ as a sum (mod $2$) of different variable conjunctions. That is, we can write $f=\sum_{i=1}^{\mathfrak{R}}\mathfrak{c}_{i}$, where each $\mathfrak{c}_{i}$ is the logical AND of a certain group of input variables. By combining variables belonging to the same party, every $\mathfrak{c}_{i}$ becomes the conjunction of at most $N$ variables, each one belonging to a different party. In this work, we restrict attention to functions $f$ that admit an ANF whose conjunctions involve no more than two variables. This covers the entire class of two-party functions, but also includes certain multi-party functions, such as the three-party majority function $\varphi_{3}(x,y,z)=xy+xz+yz\mod 2$, which outputs the majority value among inputs $x,y,z\in\{0,1\}$. In general, the functions we consider have the form $f=\sum_{i=1}^{\mathfrak{R}}a_{i}b_{i}+\sum_{k=1}^{N}z_{k}$, where $a_{i}$ and $b_{i}$ are the inputs bits of each quadratic conjunction, belonging to different parties, and $z_{k}$ is the input bit of the linear part of $f$, belonging to party $\mathbf{P}_{k}$. We let $\tilde{\mathbf{x}}_{k}\subset\{a_{i},b_{i},z_{k}\}_{i=1,k=1}^{\mathfrak{R},N}$ denote all the inputs of $f$ belonging to $\mathbf{P}_{k}$. Furthermore, if each $\tilde{\mathbf{x}}_{k}$ is no more than $M$ bits, then $\mathfrak{R}\leq\binom{N}{2}(M-1)^{2}$.
@@
The offline phase of $\mathcal{P}$ calls for the distribution of $\mathfrak{R}$ copies of the graph state |$\ket{G}$ depicted in Figure 5, one copy for each conjunction $a_{i}b_{i}$ in $f$. The generation of each copy of |$\ket{G}$ requires only two auxiliary spins in either of the experimental schemes we propose [37]. We denote the $i^{\text{th}}$ copy as |$\ket{G_{i}}$, and their distribution can be conducted in parallel. The specific qubits in each |$\ket{G_{i}}$ are given to the parties $\{\mathbf{A}_{i},\mathbf{B}_{i}\}\subset\{\mathbf{P}_{1},\cdots,\mathbf{P}_{N}\}$ who, respectively, have inputs $\{a_{i},b_{i}\}$ to the conjunction $a_{i}b_{i}$. We assume without loss of generality that each party receives qubits belonging to at least one |$\ket{G_{i}}$, an assumption that can be trivially ensured by adding conjunctions to $f$ of inputs that are identically zero. This is a technical requirement of our protocol since the |$\ket{G_{i}}$ will ultimately be used to generate one-time pad bits, and each party needs at least one. In practice, the |$\ket{G_{i}}$ can be generated by an untrusted quantum Source, and their correctness can be certified using self-testing methods (see Appendix B.3). Specific steps for building |$\ket{G}$ using a quantum emitter are presented in Appendix C.2.
@@
The online phase of $\mathcal{P}$ thereafter is split into two stages. Stage I involves measuring Pauli observables, $X,Y,Z$, on $\mathfrak{R}$ copies of the graph state |$\ket{G}$, depicted in Figure 5, following the sequence below.
@@
Input: For $i\in\{1,\cdots,\mathfrak{R}\}$ a copy of |$\ket{G_{i}}$ is distributed by the Source to $\mathbf{R}$ (the Referee) and pair of parties $\mathbf{A}_{i}$ (Alice) and $\mathbf{B}_{i}$ (Bob). The following measurement sequence is performed by $\{\mathbf{A}_{i},\mathbf{B}_{i}\}\subset\{\mathbf{P}_{1},\cdots,\mathbf{P}_{N}\}$ who holds inputs $\{a_{i},b_{i}\}$, respectively.
@@
$\beta_{i}$ $=(m_{i,5}+a_{i})b_{i}+m_{i,8}.$ (6b)
The end result of each iteration $i$ of the measurement sequence leaves Alice, Bob, and the Referee with additive homomorphic shares of the bit conjunction $a_{i}b_{i}$. The only temporal restrictions on the sequence, are that measurements in steps (I.1)-(I.3) need be performed prior to those in steps (I.4)-(I.6). Furthermore, we can understand both $\alpha_{i}$ and $\beta_{i}$ as two layers of data that are nest together by one-time pads. For $\alpha_{i}$, the bit $(m_{i,4}+m_{i,11})$ serves as a pad (known only to Alice) for the two multipliers $a_{i}$ and $(m_{i,1}+b_{i}+1)$; while within the latter multiplier, the bit $m_{i,1}$ serves as a pad (known only to Bob) for $b_{i}+1$. A similar interpretation holds for $\beta_{i}$.
@@
Figure 6: Malicious attacks on $\mathcal{P}$ by $\mathbf{S}$ and $\mathbf{R}$. (a) General attack by $\mathbf{S}$. The attack is described by a quantum instrument $\mathcal{I}$ that acts on the qubits of |$\ket{G}^{\otimes\mathfrak{R}}$ distributed to $\mathbf{S}$, with possible dependence on some side-information $\lambda.$ For every $\tilde{\mathbf{x}}$ of $\mathbf{P}$, the attack will generate for $\mathbf{S}$ a classical-quantum state $\rho_{\text{cq}}^{\mathbf{S}}(\tilde{\mathbf{x}},\lambda)$ that depends on all public communication in the protocol. We show there exists a simulator $\mathbb{S}$ in place of party $\mathbf{S}$ in the ideal world that submits some input $\tilde{\mathbf{e}}$ to $f_{\text{ideal}}$ and then uses the output $f(\tilde{\mathbf{x}},\tilde{\mathbf{e}})$ to generate the same $\rho_{\text{cq}}^{\mathbf{S}}(\tilde{\mathbf{x}},\lambda)$ achieved in the real world. (b) General attack by $\mathbf{R}$. We again show there exists a simulator $\mathbb{S}$ that takes the places of $\mathbf{R}$ in the ideal world, receiving just the function output $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ and outputting the classical-quantum state $\rho_{\text{cq}}^{\textbf{R}}(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ achieved in the real world, as shown to the right.
@@
Note that for $N$ parties $\mathbf{P}_{1},\cdots,\mathbf{P}_{N}$ these claims allow for the possibility that any subset of $N-1$ parties are colluding against a single party. The colluding parties are allowed to share an unlimited amount of classical and quantum communication over side channels, and they can thus collectively be views as a single party $\mathbf{S}$. In this case, the task essentially reduces to a problem between the two parties $\mathbf{P}$ and $\mathbf{S}$, and $\mathbf{R}$. We let $\tilde{\mathbf{x}}$ denote the input of $\mathbf{P}$ and $\tilde{\mathbf{y}}$ as the total input of all the colluding parties constituting $\mathbf{S}$. The goal, then, is to securely compute $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$.
@@
In this work we adopt a simulation based notion of security. Consider first an ideal world in which there exists some device $f_{\text{ideal}}$ that privately receives inputs $\tilde{\mathbf{x}}$ and $\tilde{\mathbf{y}}$ from $\mathbf{P}$ and $\mathbf{S}$, respectively, and then outputs $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ to $\mathbf{P}$, $\mathbf{S}$, and $\mathbf{R}$. These two worlds are described more thoroughly in Appendix B.1. Intuitively, we want the ideal world to reveal at least as much information about the inputs $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ than what is revealed to any dishonest party acting maliciously in the real world.
@@
This intuition is made precise through the notion of a simulator. A simulator, $\mathbb{S}$, in this setting is some device that can take the place of a party $\mathbf{X}\in\{\mathbf{P},\mathbf{S},\mathbf{R}\}$ and interact with the ideal functionality $f_{\text{ideal}}$ from the viewpoint of party $\mathbf{X}$. Consider now any action taken by a potentially dishonest party in $\mathcal{P}$. Our security definition is that there must exists a simulator $\mathbb{S}$ interacting with $f_{\text{ideal}}$ that exactly reproduces what the dishonest party obtains from $\mathcal{P}$. This operationally captures the idea that the adversary learns no more information about an honest party’s input, other than what can be inferred directly from an ideal evaluation of $f$. Figure 6 depicts the two forms of malicious attacks for which we prove that $\mathcal{P}$ is secure. A formal security definition and full proof are provided in Appendix B.1.
@@
Performance. To handle experimental errors in the above protocol, we can employ a simple repetition code to suppress the effects of any infidelity in our ability to make each copy of |$\ket{G}$ in the offline phase of $\mathcal{P}$. By determining the total bit error probability associated with each share of $a_{i}b_{i}$, an arbitrarily small total bit error probability $\epsilon_{f}$ on the output $f$ can be chosen to set the number of repetitions required for Stage I. We use this information to estimate the lower bound rate of computation at which $N$ parties can compute $f$ on their $M$-bit inputs. Details are provided in Appendix B.2.
@@
Figure 7: Error corrected two-party computation rate versus the number of requiring conjunctions in $f$ using the HoD scheme, shown for two acceptable error probabilities on the computation ($\epsilon_{f}$). The rate is expressed in units of the repetition rate of excitation of the chosen quantum emitter, $R_{\text{rep}}$, which can be in the range of $10^{6}-10^{9}~{}\rm{s}^{-1}$ for highly coherent emitters. The individual bit error probabilities for each copy of |$\ket{G}$ is 0.012, assuming $\eta_{e}=0.4$, $F_{\text{ent}}=0.999$, and $t_{\text{rep}}/\tau=10^{-9}$ (see. Appendix D for details). We see that these parameters allow virtually unlimited reduction in the total error probability with minimal change in the overall rate.

### v3

@@ II Emit-then-add
where we include a superscript to denote the emitter $(e)$ and auxiliary spins $(s)$, respectively. For simplicity in the above, we have used the same coherence time $\tau$ for the emitter and auxiliary spins in the system, though this may not be the case for hybrid systems [28, 27].
[…]
Scaling. Finally, the most significant advantage of emit-then-add comes in scaling up the size of graph states using state-of-the-art to near-term hardware. The average time to successfully generate a $n_{p}$-photon graph via the deterministic scheme, by directly collecting photons from an emitter in $n_{p}$ subsequent excitation events, is $\mathcal{O}(\eta_{e}^{-n_{p}})$, where $\eta_{e}$ is the emitter collection efficiency, because any failed detection event truncates the graph. For our emit-then-add scheme, a failed detection of the emitter photon simply triggers reinitialization of the emitter and another attempt at photon emission, while the graph under construction remains unaffected. Therefore, a $n_{p}$-photon graph state will be created over a time that is $\mathcal{O}(n_{p}\eta_{e}^{-1})$.
As a comparison, the time to make a $n_{p}$-photon graph state (in units of the repetition period for exciting the emitter) is shown in Fig. 2 for $\eta_{e}\in\{0.1,0.5,0.9\}$ for deterministic (red), and emit-then-add (blue) schemes. We do not include the time to perform gates and MCMR operations of the spins for these estimates, and we assume that the emitter coherence does not limit the graph size for either scheme (i.e. $\tau/n_{p}\gg t_{\text{rep}}/\eta_{e}+t_{\text{add}}$). We include cutoffs where the additional infidelity, captured by $F_{\text{add}}$ defined above, precludes the construction of larger graph states. Explicitly, given the current state-of-the-art for trapped ion and neutral systems [45, 46, 47, 48, 49], the vertical gray lines in Fig. 2 represent where the fidelity of the final graph state generated with emit-then-add falls below $1/2$.
III Two-party computation with graph states
@@
As one application of our emit-then-add scheme, we describe a method for securely computing an arbitrary Boolean function $f:\{0,1\}^{\times n}\to\{0,1\}$ of either two parties or a restricted class of multi-party functions. As a special type of MBQC, our protocol, $\mathcal{P}$, performs the calculation through a sequence of measurements on distributed graph states, followed by classical broadcasting and local processing. Only Pauli measurements are needed, so $\mathcal{P}$ can be implemented without the need for photonic memory, where the generation of the requisite resource state happens in parallel with the computation. Furthermore, the size of each graph state required for $\mathcal{P}$ is fixed at $12$ photons, regardless of the number of parties or the size of their inputs, which bodes well for state-of-the-art to near-term hardware.
@@
We propose a method for restricted MPC that includes all two-party computations and requires only two rounds of public communication in the form broadcasting, or openings, regardless of the size of the computational input. Our protocol follows a well-known approach of decomposing an MPC into offline and online phases [53, 54, 55]. In the offline phase, some universal computational resource is distributed to all the parties. Crucially, this resource does not depend on the particular function being computed other than its input size. Then in the online phase this resource is used to compute some chosen function of the parties’ inputs. For example, in the classical setting one well-known computational resource is a special form of shared randomness known as “Beaver triples,” which can be used to efficiently compute logical AND gates in the online phase [56]. The problem of MPC then reduces in part to secure and efficient offline methods for distributing shared randomness, such as Beaver triples, which directly enable secure online computation. In a similar spirit, our protocol involves distributing certain quantum graph states in the offline phase which then enable the computation of a logical AND in the online phase, through quantum measurement and classical post-processing.
Beyond its relatively low communication costs, a significant advantage of our protocol is that the parties can, in principle, use self-testing methods to unconditionally verify that some untrusted Source is faithfully distributing the correct graph state [57, 58], an ability that does not exist for classical sources of shared randomness. Furthermore, we prove that our online phase offers unconditional privacy against an arbitrary malicious adversary, with access to a general quantum instrument [59], in the sense that honest participation reveals no information about a party’s input, other than what can be inferred from the evaluation of the final computed function.
Note that in our discussion of the protocol here we do not offer any means of ensuring fairness, such security with abort or guaranteed output delivery. We do not claim that the full MPC here is unconditionally secure in this sense, and a classical means of ensuring fairness for the broadcasts in our online phase can be employed with computational security guarantees [60, 61]. Instead, the novelty of this work is in the ability to disseminate a specific online correlation with unconditional security. In a similar vein as quantum key distribution, we demonstrate how quantum states offer the ability to distribute information-theoretic secure shared randomness, in our case, in a form that directly enables MPC.
@@
Overview. Suppose that $N$ parties $\mathbf{P}_{1},\cdots,\mathbf{P}_{N}$ wish to compute some Boolean function $f(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$, where $\mathbf{x}_{k}$ is a string of bits representing the input for party $\mathbf{P}_{k}$. In addition to correctness, the evaluation of $f$ should be done securely such that the parties learn no more information about the individual $\mathbf{x}_{1},\cdots,\mathbf{x}_{N}$ beyond their own input and what is revealed in the function value $f(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$. To achieve this task, we propose a method of delegated computation in which a non-collaborating Referee, $\mathbf{R}$, is introduced to assist in the computation of $f(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$. To maintain privacy, $\mathbf{R}$ also should not learn any more information about the $\mathbf{x}_{k}$ beyond what is implied by the computed value $f(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$, nor does $\mathbf{R}$ reveal any more information to the other parties, other than what is necessary to evaluate $f$ securely.
@@
We utilize the fact that every Boolean function $f$ can be expressed in an algebraic normal form (ANF), which presents $f$ as a sum (mod $2$) of different variable conjunctions. That is, we can write $f=\sum_{i=1}^{\mathfrak{R}}\mathfrak{c}_{i}$, where each $\mathfrak{c}_{i}$ is the logical AND of a certain group of input variables. By combining variables belonging to the same party, every $\mathfrak{c}_{i}$ becomes the conjunction of at most $N$ variables, each one belonging to a different party. In this work, we restrict attention to functions $f$ that admit an ANF whose conjunctions involve no more than two variables. This covers the entire class of two-party functions, but also includes certain multi-party functions, such as the three-party majority function $\varphi_{3}(x,y,z)=xy+xz+yz\mod 2$, which outputs the majority value among inputs $x,y,z\in\{0,1\}$. In general, the functions we consider have the form $f=\sum_{i=1}^{\mathfrak{R}}a_{i}b_{i}+\sum_{k=1}^{N}z_{k}$, where $a_{i}$ and $b_{i}$ are the input bits of each quadratic conjunction, belonging to different parties, and $z_{k}$ is the input bit of the linear part of $f$, belonging to party $\mathbf{P}_{k}$. We let $\tilde{\mathbf{x}}_{k}\subset\{a_{i},b_{i},z_{k}\}_{i=1,k=1}^{\mathfrak{R},N}$ denote all the inputs of $f$ belonging to $\mathbf{P}_{k}$. Furthermore, if each $\tilde{\mathbf{x}}_{k}$ is no more than $M$ bits, then $\mathfrak{R}\leq\binom{N}{2}(M-1)^{2}$.
@@
The offline phase of $\mathcal{P}$ calls for the distribution of $\mathfrak{R}$ copies of the graph state |$\ket{G}$ depicted in Fig. 3, one copy for each conjunction $a_{i}b_{i}$ in $f$. The generation of each copy of |$\ket{G}$ requires only two auxiliary spins in the experimental schemes we propose [40]. We denote the $i^{\text{th}}$ copy as |$\ket{G_{i}}$, and their distribution can be conducted in parallel. The specific qubits in each |$\ket{G_{i}}$ are given to the parties $\{\mathbf{A}_{i},\mathbf{B}_{i}\}\subset\{\mathbf{P}_{1},\cdots,\mathbf{P}_{N}\}$ who, respectively, have inputs $\{a_{i},b_{i}\}$ to the conjunction $a_{i}b_{i}$. We assume without loss of generality that each party receives qubits belonging to at least one |$\ket{G_{i}}$, an assumption that can be trivially ensured by adding conjunctions to $f$ of inputs that are identically zero. This is a technical requirement of our protocol since the |$\ket{G_{i}}$ will ultimately be used to generate one-time pad bits, and each party needs at least one. In practice, the |$\ket{G_{i}}$ can be generated by an untrusted quantum Source, and their correctness can be certified using self-testing methods (see Appendix D.3). Specific steps for building |$\ket{G}$ in the experimental schemes given in Sec. II are presented in Appendix B.2.
@@
The online phase of $\mathcal{P}$ thereafter is split into two stages. Stage I involves measuring Pauli observables, $X,Y,Z$, on $\mathfrak{R}$ copies of the graph state |$\ket{G}$, depicted in Fig. 3, following the sequence below.
@@
Input: For $i\in\{1,\cdots,\mathfrak{R}\}$, the following measurement sequence is performed by $\{\mathbf{A}_{i},\mathbf{B}_{i}\}\subset\{\mathbf{P}_{1},\cdots,\mathbf{P}_{N}\}$, along with $\mathbf{R}$, on copy |$\ket{G_{i}}$. $\mathbf{A}$ inputs $a_{i}$.
@@
$\beta_{i}$ $=(m_{i,5}+a_{i})s_{i}+m_{i,8}.$ (4b)
The end result of each iteration $i$ of the measurement sequence leaves Alice, Bob, and the Referee with additive homomorphic shares of the bit conjunction $a_{i}b_{i}$. The only temporal restrictions on the sequence, are that measurements in steps (I.1)-(I.3) need be performed prior to those in steps (I.4)-(I.6). Furthermore, we can understand both $\alpha_{i}$ and $\beta_{i}$ as two layers of data that are nested together by one-time pads. For $\alpha_{i}$, the bit $(m_{i,4}+m_{i,11})$ serves as a pad (known only to Alice) for the two multipliers $a_{i}$ and $(m_{i,1}+b_{i}+1)$; while within the latter multiplier, the bit $m_{i,1}$ serves as a pad (known only to Bob) for $b_{i}+1$. A similar interpretation holds for $\beta_{i}$.
@@
Figure 4: Malicious attacks on $\mathcal{P}$ by $\mathbf{S}$ and $\mathbf{R}$. Real (left) vs. ideal (right) instantiations of the attacks are depicted, where boxes (red, blue, green) indicate local processing by a given party ($\mathbf{P},\mathbf{S},\mathbf{R}$). (a) General attack by $\mathbf{S}$. The attack is described by a quantum instrument $\mathcal{I}$ that acts on the qubits of |$\ket{G}^{\otimes\mathfrak{R}}$ distributed to $\mathbf{S}$, with possible dependence on some side-information $\lambda.$ For every $\tilde{\mathbf{x}}$ of $\mathbf{P}$, the attack will generate for $\mathbf{S}$ a classical-quantum state $\rho_{\text{cq}}^{\mathbf{S}}(\tilde{\mathbf{x}},\lambda)$ that depends on all public communication in the protocol. We show there exists a simulator $\mathbb{S}$ in place of party $\mathbf{S}$ in the ideal world that submits some input $\tilde{\mathbf{e}}$ to $f_{\text{ideal}}$ and then uses the output $f(\tilde{\mathbf{x}},\tilde{\mathbf{e}})$ to generate the same $\rho_{\text{cq}}^{\mathbf{S}}(\tilde{\mathbf{x}},\lambda)$ achieved in the real world. (b) General attack by $\mathbf{R}$. We again show there exists a simulator $\mathbb{S}$ that takes the places of $\mathbf{R}$ in the ideal world, receiving just the function output $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ and outputting the classical-quantum state $\rho_{\text{cq}}^{\textbf{R}}(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ achieved in the real world, as shown to the right.
@@
Note that for $N$ parties $\mathbf{P}_{1},\cdots,\mathbf{P}_{N}$ these claims allow for the possibility that any subset of $N-1$ parties are colluding against a single party. The colluding parties are allowed to share an unlimited amount of classical and quantum communication over side channels, and they can thus collectively be viewed as a single party $\mathbf{S}$. In this case, the task essentially reduces to a problem between the two parties $\mathbf{P}$ and $\mathbf{S}$, and $\mathbf{R}$. We let $\tilde{\mathbf{x}}$ denote the input of $\mathbf{P}$ and $\tilde{\mathbf{y}}$ as the total input of all the colluding parties constituting $\mathbf{S}$. The goal, then, is to securely compute $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ with an honest majority in $\{\mathbf{P},\mathbf{S},\mathbf{R}\}$.
@@
In this work we adopt a simulation based notion of security. Consider first an ideal world in which there exists some device $f_{\text{ideal}}$ that privately receives inputs $\tilde{\mathbf{x}}$ and $\tilde{\mathbf{y}}$ from $\mathbf{P}$ and $\mathbf{S}$, respectively, and then outputs $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ to $\mathbf{P}$, $\mathbf{S}$, and $\mathbf{R}$. These two worlds are described more thoroughly in Appendix D.1. Intuitively, we want the ideal world to reveal at least as much information about the inputs $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ than what is revealed to any dishonest party acting maliciously in the real world.
@@
This intuition is made precise through the notion of a simulator. A simulator, $\mathbb{S}$, in this setting is some device that can take the place of a party $\mathbf{X}\in\{\mathbf{P},\mathbf{S},\mathbf{R}\}$ and interact with the ideal functionality $f_{\text{ideal}}$ from the viewpoint of party $\mathbf{X}$. Consider now any action taken by a potentially dishonest party in $\mathcal{P}$. Our security definition is that there must exists a simulator $\mathbb{S}$ interacting with $f_{\text{ideal}}$ that exactly reproduces what the dishonest party obtains from $\mathcal{P}$. This operationally captures the idea that the adversary learns no more information about an honest party’s input, other than what can be inferred directly from an ideal evaluation of $f$. Fig. 4 depicts the two forms of malicious attacks for which we prove that $\mathcal{P}$ is secure. A formal security definition and full proof are provided in Appendix D.1.
@@
Performance. To handle experimental errors in the above protocol, we can employ a simple repetition code to suppress the effects of any infidelity in our ability to make each copy of |$\ket{G}$ in the offline phase of $\mathcal{P}$. By determining the total bit error probability associated with each share of $a_{i}b_{i}$, an arbitrarily small total bit error probability $\epsilon_{f}$ on the output $f$ can be chosen to set the number of repetitions required for Stage I. We use this information to estimate the lower bound rate of computation at which $N$ parties can compute $f$ on their $M$-bit inputs. Details are provided in Appendix D.2.
@@
Figure 5: Error corrected two-party computation rate versus the number of required conjunctions in $f$, shown for two acceptable error probabilities on the computation ($\epsilon_{f}$). We assume the use the our scheme with destructive measurements and no photonic memory. The rate is expressed in units of the repetition rate of excitation of the chosen quantum emitter, $t_{\text{rep}}^{-1}$, which can be in the range of $10^{6}-10^{9}~{}\rm{s}^{-1}$ for highly coherent emitters. The individual bit error probability for each copy of |$\ket{G}$ is a pessimistic 0.157, assuming $\eta_{e}=0.1$, and $F_{\text{add}}=0.99$ (see. Appendix C and D.2 for details). We see that even these parameters allow virtually unlimited reduction in the total error probability with minimal change in the overall rate.